# EgoExo-Fitness: Towards Egocentric and Exocentric Full-Body Action Understanding

### v1

@@ 38 SB, CJ, LRL   76 SU, SB, LRL, KRAMC, CJ, SS
Implementation. (1) Data Construction: We select 4,753 single action videos (3,000 for training and 1,753 for testing) to construct the Action Classification benchmark. (2) Pre-trained weights: We evaluate models with various pre-training strategies to construct action classification benchmark. For I3D [9], EgoVLP [42], and TimeSformer [4] pretrained on the K600 [8] dataset, we use the official pretrained weights. For TimeSformer pre-trained on Ego-Exo4D [23], we follow the setting of “Key-Step Recognition” benchmark in Ego-Exo4D to initialize the model with K600 pre-trained weights then trained on Ego-Exo4D. (3) Experiment Settings: The input size of the video clip is set as $16\times 224\times 224$. During training, the video clips are sampled with temporal augmentation followed by random cropping. We train the models for 200 epochs with a base learning rate of 1e-5 and adopt a multi-step learning rate decay with a decay rate of 0.5 for every 25 epochs. For evaluation, a single video is uniformly sampled from the video, followed by center cropping.
@@
Specifically, in training phase, a training set $D_{train}=\{(v_{i},s_{i})\}^{N}_{i=1}$ is used to construct a sequence classification task, where $v_{i}$ is a action sequence video and $s_{i}$ is a sequence label (e.g., a SID in Tab. 1). Given a video $v\in\mathbb{R}^{3\times H\times W\times T}$ and its corresponding sequence label $s$, the model $f\odot g:{R}^{3\times H\times W\times T}\rightarrow\mathbb{R}^{C}$ is asked to predict the sequence label from $C$ sequence classes. Here $f$ is the embedding encoder, $g$ is the classifier. $H$, $W$, and $T$ are height, width, and the number of frames, respectively. In the testing phase, the model is asked to perform sequence verification on the test set where the sequence labels do not overlap with videos in the training set. Given a video pair $(v_{i},v_{j})$, a distance (or similarity) function $D$ is conducted on the embeddings of each video in the pair, which is denoted as $d_{ij}=D(f(v_{i}),f(v_{j}))$. A higher $d_{ij}$ indicates a lower possibility for $v_{i}$ to contain the same action sequence as $v_{j}$ (opposite if similarity function is used). In practical application, a threshold $\tau$ can be set to decide whether two sequences are consistent: if $d_{ij}>\tau$, sequences of $v_{i}$ and $v_{j}$ are consistent, otherwise inconsistent (opposite if similarity function is used).
More Details about baseline model. As discussed in the main paper, we adopt the state-of-the-art sequence verification model CAT [53] as the baseline model. The overview of CAT is shown in Fig. A.4. The embedding encoder $f$ includes a 2D Backbone and a Temporal Modeling Module to encode video embeddings. The classifier $g$ is implemented as a Multi-Layer Perceptron. Specifically, the 2D backbone is implemented as a CLIP-ViT/16 [54], and the Transformer encoder is adopted as a Temporal Modeling Module. During training, except for the classification loss $L_{CLS}$, an extra sequence alignment loss $L_{SA}$ is adopted to align video representations of videos with the same action sequence. For more details about the loss function, please refer to the original works [53, 14].
@@
Table A.7: More comparison betweens the proposed EgoExo-Fitness and the concurrent EgoExo4D [23] dataset. Compared with EgoExo4D, the proposed EgoExo-Fitness collects videos of a new scenario (i.e., fitness) and augments data with novel annotations of interpretable action judgment (i.e., text guidance and technical keypoint verification are not provided in EgoExo4D).
Datasets Scenarios Step Text Keypoint Comments Score Durations
@@
EgoExo-Fitness(Ours) Fitness 31h
@@
# Videos 1412 303 3000 7200 48 18 1248
# Single Actions 1412 32697 3000 7200 - 615 6211

### v2

@@ 29 KPU, SU, KTT, LLKL, JJ, HK 58 SB, CJ, KRAMC, KPU, SS
Implementation. (1) Data Construction: We select 4,753 single action videos (3,000 for training and 1,753 for testing) to construct the Action Classification benchmark. (2) Pre-trained weights: We evaluate models with various pre-training strategies to construct action classification benchmark. For I3D [9], EgoVLP [43], and TimeSformer [4] pre-trained on the K600 [8] dataset, we use the official pre-trained weights. For TimeSformer pre-trained on Ego-Exo4D [24], we follow the setting of “Key-Step Recognition” benchmark in Ego-Exo4D to initialize the model with K600 pre-trained weights then trained on Ego-Exo4D. (3) Experiment Settings: The input size of the video clip is set as $16\times 224\times 224$. During training, the video clips are sampled with temporal augmentation followed by random cropping. We train the models for 200 epochs with a base learning rate of 1e-5 and adopt a multi-step learning rate decay with a decay rate of 0.5 for every 25 epochs. For evaluation, a single video is uniformly sampled from the video, followed by center cropping.
@@
Specifically, in training phase, a training set $D_{train}=\{(v_{i},s_{i})\}^{N}_{i=1}$ is used to construct a sequence classification task, where $v_{i}$ is a action sequence video and $s_{i}$ is a sequence label (e.g., a SID in Tab. A.2). Given a video $v\in\mathbb{R}^{3\times H\times W\times T}$ and its corresponding sequence label $s$, the model $f\odot g:{R}^{3\times H\times W\times T}\rightarrow\mathbb{R}^{C}$ is asked to predict the sequence label from $C$ sequence classes. Here $f$ is the embedding encoder, $g$ is the classifier. $H$, $W$, and $T$ are height, width, and the number of frames, respectively. In the testing phase, the model is asked to perform sequence verification on the test set where the sequence labels do not overlap with videos in the training set. Given a video pair $(v_{i},v_{j})$, a distance (or similarity) function $D$ is conducted on the embeddings of each video in the pair, which is denoted as $d_{ij}=D(f(v_{i}),f(v_{j}))$. A higher $d_{ij}$ indicates a lower possibility for $v_{i}$ to contain the same action sequence as $v_{j}$ (opposite if similarity function is used). In practical application, a threshold $\tau$ can be set to decide whether two sequences are consistent: if $d_{ij}>\tau$, sequences of $v_{i}$ and $v_{j}$ are consistent, otherwise inconsistent (opposite if similarity function is used).
More Details about baseline model. As discussed in the main paper, we adopt the state-of-the-art sequence verification model CAT [55] as the baseline model. The overview of CAT is shown in Fig. A.4. The embedding encoder $f$ includes a 2D Backbone and a Temporal Modeling Module to encode video embeddings. The classifier $g$ is implemented as a Multi-Layer Perceptron. Specifically, the 2D backbone is implemented as a CLIP-ViT/16 [56], and the Transformer encoder is adopted as a Temporal Modeling Module. During training, CAT takes a pair of videos with the same action sequence as input and is optimized to learn to classify the action sequence labels by a classification loss $L_{CLS}$. Besides, an extra sequence alignment loss $L_{SA}$ is adopted to align video representations of videos with the same action sequence.
@@
Table A.8: More comparison betweens the proposed EgoExo-Fitness and the concurrent Ego-Exo4D [24] dataset. Compared with Ego-Exo4D, the proposed EgoExo-Fitness collects videos of a new scenario (i.e., fitness) and augments data with novel annotations of interpretable action judgment (i.e., text guidance and technical keypoint verification are not provided in Ego-Exo4D).
Datasets Scenarios Step Text Keypoint Comment Score Duration
@@
EgoExo-Fitness(Ours) Fitness 32h
@@
# Videos 1412 303 3000 7200 48 18 1276
# Single Actions 1412 32697 3000 7200 - 615 6131